(b)  

(a) Write down the equation for a sinusoidal voltage of 50 Hz and its peak value is 20 V. 

phase angle between voltage and current. 


1
Expert's answer
2021-10-07T03:15:31-0400

(a) The equation is


"v(t)=V_\text{max}\sin(2\pi ft+\phi),\\
v(t)=20\sin(314t+\phi)."